KTeV shift Summary, Day shift 12/11/1999

Tony Barker (tonyb@cuhep.Colorado.EDU)
Fri, 10 Dec 1999 15:44:16 -0700 (MST)

Shift Summary for Day Shift, 11 December 1999

On Shift: T.Barker (J.Ladue from 13:00 to 15:00)

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7.5

Accounting of lost time: Tev STILL down.
Estimate is now Monday for beam.
========================================================================
Runs ended this shift: none
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: Surveying crew; RSK to work on KAMI readout
========================================================================
Overview of shift: J.Shields continues skim work.
Hogan turned chambers off to facilitate survey
work; he will turn them on again during the
next shift. New gain loaded for CsI channel
1096 which received a cookie-transplant recently.
=======================================================================
Follow up Notes on Previous Shift
none
========================================================================
Items to Follow up on Next Shift
Chambers should be turned on again.